# TiCoSS: Tightening the Coupling between Semantic Segmentation and Stereo Matching within A Joint Learning Framework

### v3

@@ Abstract
Semantic segmentation and stereo matching, respectively analogous to the ventral and dorsal streams in our human brain, are two key components of autonomous driving perception systems. Addressing these two tasks with separate networks is no longer the mainstream direction in developing computer vision algorithms, particularly with the recent advances in large vision models and embodied artificial intelligence. The trend is shifting towards combining them within a joint learning framework, especially emphasizing feature sharing between the two tasks. The major contributions of this study lie in comprehensively tightening the coupling between semantic segmentation and stereo matching. Specifically, this study introduces three key novelties: (1) a tightly coupled, gated feature fusion strategy, (2) a hierarchical deep supervision strategy, and (3) a coupling tightening loss function. All these technical contributions help tightly couple semantic segmentation and stereo matching tasks, collectively resulting in TiCoSS, a state-of-the-art joint learning framework. Through extensive experiments on the KITTI and vKITTI2 datasets, along with both qualitative and quantitative analyses, we validate the effectiveness of our developed strategies and loss function. Our approach demonstrates superior performance compared to prior arts, with a notable increse in mIoU by over 9%. Our source code will be publicly available at mias.group/TiCoSS upon publication.
@@
semantic segmentation, stereo matching, joint learning, computer vision, artificial intelligence.
@@ I-A Background
VISUAL environment perception (VEP) serves as a fundamental and front-end module in robotic systems [1]. Semantic segmentation and stereo matching are two key VEP tasks [2]. The former, akin to the ventral stream in our brain, provides a pixel-level understanding of the scene, while the latter, akin to the dorsal stream in our brain, mimics human binocular vision to acquire accurate depth information, essential for 3D geometry reconstruction [3]. These two tasks collaborate to deliver both contextual and geometric information, resulting in a comprehensive scene understanding that significantly enhances the capabilities of robotic systems [4].
Nevertheless, previous studies [5, 6, 7, 8, 9] address these two tasks with separate networks, which limits their potential to share informative contextual and geometric information [10]. For instance, stereo matching networks can occasionally produce ambiguous disparity estimations, particularly in texture-less and occluded regions [11]. Semantic segmentation can provide pixel-level scene understanding to disambiguate these cases, ultimately leading to more reliable disparity estimations [12]. In addition, semantic segmentation networks often struggle to distinguish clear object boundaries, particularly in complex driving scenarios, due to the lack of spatial geometric information [13]. A common solution for improved semantic segmentation performance is to employ feature-fusion networks equipped with duplex encoders to extract heterogeneous features from RGB-X data [14, 15, 16], where “X” provides spatial geometric information, such as the depth images generated from LiDAR point clouds [17] and surface normal maps obtained through depth-to-normal translation [18]. However, the availability and quality of “X” significantly influence the overall performance of semantic segmentation and can potentially limit the widespread adoption of feature-fusion networks [19].
Therefore, in recent years, the simultaneous learning, deployment, and inference of both tasks have become a mainstream [20, 21, 16, 22, 23], because a unified joint learning framework can process contextual and geometric information more comprehensively. It also enables end-to-end training of the entire system, capable of tackling the challenges posed by both tasks [16]. Consequently, this can enhance the overall performance in both semantic segmentation and stereo matching tasks, outperforming models trained separately for each task [23]. Moreover, a unified framework can potentially reduce computational complexity and memory usage, as learning representations can likely be shared between both tasks [24]. This can be advantageous particularly for resource-constrained applications that require real-time performance.
@@ I-B Existing Challenges and Motivation
The performance of a feature-fusion semantic segmentation network is heavily influenced by the employed strategy for heterogeneous feature fusion [25, 26]. Currently, the bottleneck lies in the simplistic and indiscriminate fusion of heterogeneous features, which often causes conflicting learning representations and erroneous segmentation results [27]. Taking our previous work S3M-Net [16] as an example, its adopted feature fusion strategy essentially performs an element-wise summation between contextual and geometric feature maps at each stage. These feature maps are then directly fed into subsequent layers without excluding irrelevant information, leading to a loose coupling between the semantic segmentation and stereo matching tasks within the encoder. Furthermore, as the network goes deeper, such an indiscriminate feature fusion strategy tends to diminish the proportion of useful geometric features in the decoder’s input [28], potentially leading to unsatisfactory semantic segmentation performance.
Additionally, due to the vanishing gradient problem [29], existing joint learning frameworks usually exhibit slow convergence during training. A common solution is to employ the deep supervision (DS) strategies that incorporate additional pathways to achieve gradient propagation [30]. Nonetheless, existing DS strategies employed in feature-fusion semantic segmentation networks typically overlook the potential interactions between the main and side auxiliary classifiers, failing to exploit the complementarity among all classifiers and can therefore restrict the improvements in semantic segmentation performance within our multi-task joint learning framework.
In the loss function aspect, previous joint learning frameworks, such as SegStereo[21] and SSNet [31], typically compute the losses for two tasks separately and supervise the entire training process by simply minimizing the weighted sum of these losses. Such loss function fails to exploit the complementarity between these two tasks at the output level, also resulting in their loose coupling.
Prior joint learning arts, such as S3M-Net [16] and DSNet [10], primarily focus on introducing a joint learning framework that performs semantic segmentation and stereo matching simultaneously. However, exploring the potential complementarity and tightening the coupling between these two tasks, particularly in the encoder, decoder and loss function aspects, have received relatively limited attention in this research area and calls for further investigation.
@@ I-C Contributions
To address the aforementioned limitations, we introduce Tightly-Coupled Semantic Segmentation and Stereo Matching Network (TiCoSS), an end-to-end joint learning approach that focuses primarily on improving the coupling between stereo matching and semantic segmentation tasks, which has not been emphasized in previous related studies. Our proposed TiCoSS introduces three new techniques: (1) a tightly-coupled, gated feature fusion (TGF) strategy, which utilizes a series of selective inheritance gates (SIGs) to propagate useful contextual and geometric information from the preceding layer to the current layer; (2) a hierarchical deep supervision (HDS) strategy that uses the highest resolution fused feature maps to guide deep supervision throughout each branch, as these features contain the most abundant local spatial details; (3) a novel coupling tightening (CT) loss, consisting of a widely used stereo matching loss presented in [9], the semantic consistency-guided (SCG) loss in [16], a disparity inconsistency-aware (DIA) loss that utilizes disparity estimations to refine segmentation results, and a deep supervision consistency constraint (DSCC) loss which employs the Kullback-Leibler (KL) divergence to improve prediction consistency across outputs from all deep supervision branches. These contributions collectively advance S3M-Net, resulting in TiCoSS, a new, powerful, and tightly-coupled joint learning framework that simultaneously performs robust and accurate semantic segmentation and stereo matching tasks. Extensive experiments conducted on the vKITTI2 [32] and KITTI 2015 [33] datasets unequivocally demonstrate the effectiveness of the aforementioned contributions and the superior performance of TiCoSS over other state-of-the-art (SoTA) approaches.
@@
  • The DIA loss and the DSCC loss that tighten the coupling between the two tasks, thereby further improving semantic segmentation performance.
@@ II-A Semantic Segmentation
Semantic segmentation has been a long-standing challenge in the fields of computer vision and robotics over the past decade [3, 34]. SoTA networks generally fall into two groups [16]: (1) single-modal networks (with a single encoder) and (2) feature-fusion networks (with multiple encoders) [15, 35, 36]. Early efforts primarily focused on encoder-decoder architectures for pixel-level classification. Representative examples include PSPNet [5], the DeepLab series [6, 37], as well as Transformer-based networks [38, 7, 39]. The encoder extracts hierarchical, contextual feature maps from input images, while the decoder generates segmentation maps by upsampling and combining feature maps from different encoder layers. Nonetheless, these networks are limited in effectively combining heterogeneous features extracted from different sources (or modalities) of visual information, which makes it challenging to produce accurate segmentation results in scenarios with poor lighting and illumination conditions [15]. This has led researchers to focus on feature-fusion networks that can effectively fuse heterogeneous features learned from multiple sources (or modalities) of visual information. This problem is often referred to as “RGB-X semantic segmentation”, where “X” represents the additional modality (or source) of visual information, in addition to the RGB images. The most representative feature-fusion networks include convolutional neural network (CNN)-based ones, such as MFNet [40], RTFNet [41], SNE-RoadSeg series [15, 42, 28], and S3M-Net [16], as well as Transformer-based ones, such as OFF-Net [43] and RoadFormer [14]. In this article, we design TiCoSS based on our previously proposed S3M-Net, with a specific emphasis on exploring more effective solutions for tighter coupling between semantic segmentation and stereo matching.
@@ II-B Stereo Matching
Owing to recent advancements in deep learning techniques, end-to-end deep stereo networks [8, 44, 45, 46, 9, 47] have dramatically outperformed traditional explicit programming-based stereo matching algorithms. PSMNet [8] uses a spatial pyramid to capture multi-scale information and uses multiple 3D convolutional layers to exploit both local and global contexts for cost computation. GwcNet [44] builds on PSMNet by constructing the cost volume through group-wise correlation, enhancing the 3D stacked hourglass network. To address the computational demands of 3D convolutions, researchers have sought ways to balance efficiency and accuracy in stereo matching. LEA-Stereo [46], for instance, introduces the first neural architecture search (NAS) [48] framework to stereo matching. RAFT-Stereo [9], a rectified stereo matching approach, developed based on RAFT [49], uses a series of gated recurrent units to refine correlation features and enhance disparity estimation accuracy. CRE-Stereo [47] leverages recurrent refinement to update disparities in a coarse-to-fine fashion and adaptive group correlation to minimize the impact of erroneous rectification, leading to more compelling disparity estimation results. In this article, we primarily focus on improving semantic segmentation performance, and thus, adopt the stereo matching approach used in our previous work, S3M-Net.
@@ II-C Simultaneous Semantic Segmentation and Stereo Matching
Existing joint learning frameworks that simultaneously address these two tasks mainly focus on improving disparity accuracy by leveraging semantic information [21, 20, 12, 22, 23, 10]. However, discussions on utilizing disparity information to enhance semantic segmentation at the feature level for joint learning remain limited, except for the aforementioned “RGB-X semantic segmentation”. These prior arts often require extensive annotated training data or involve complex training strategies for joint learning. For example, SegStereo [21] necessitates an initial unsupervised training phase on the large-scale Cityscapes [50] dataset, followed by a subsequent supervised fine-tuning on the smaller KITTI [51, 33] datasets. Similarly, the approaches introduced in [12, 22, 23] require pre-training their spatial branches for stereo matching on the large-scale SceneFlow [52] dataset before fine-tuning both semantic and spatial branches on the KITTI [51, 33] datasets. DSNet [10] adopts a different joint learning strategy by alternating the training of the semantic segmentation and stereo matching networks, with parameters of each network being frozen during the training of the other. Nevertheless, leveraging both contextual and geometric information can be challenging, as the shared features are not learned in an end-to-end manner. DispSegNet [20] utilizes an embedding learned from the semantic segmentation branch to refine the initial disparity estimation. RTS2Net [22] relies on coarse-to-fine estimations in a multi-stage fashion for accurate disparity estimation. However, both methods only achieve limited improvement in segmentation results since they fail to leverage the informative spatial details to enhance semantic segmentation. Recently, SSNet [31] uses a single encoder to extract shareable features for both tasks. However, as demonstrated in [53], such shareable features may not be suitable for both dense prediction and geometric vision tasks. S3M-Net uses two separate encoding branches to achieve these two tasks simultaneously, but the weak coupling between these branches limits the integration of contextual and geometric information. Additionally, the source code (in PyTorch or TensorFlow) for these previous approaches is not publicly available, and re-implementing them carries the risk of introducing errors. In contrast to the aforementioned approaches, our proposed TiCoSS uses a tightly-coupled joint learning framework that effectively leverages both contextual and geometric information. Moreover, TiCoSS is trained in an end-to-end manner and is capable of learning accurate and robust semantic segmentation and stereo matching tasks simultaneously, even with limited training data.
@@ III-B Tightly-coupled Gated Fusion Strategy
S3M-Net [16] proposes an effective joint learning framework to simultaneously perform semantic segmentation and stereo matching. Despite achieving impressive results, these two tasks are loosely coupled. It merely employs the feature fusion strategy proposed in SNE-RoadSeg [15], where the geometric features extracted from disparity maps are indiscriminately fused into the contextual features extracted from RGB images via simplistic element-wise summation. The fused heterogeneous features are then treated as preceding contextual features and fed into subsequent layers without selective processing, which misleads the semantic segmentation task. This is primarily because the deeper geometric features contain irrelevant semantic information, and as the network goes deeper, the proportion of RGB features in the decoder’s input tends to diminish [28].
Our TGF strategy is, therefore, designed to overcome this limitation by selectively complementing contextual features with useful geometric features, resulting in a tightly-coupled duplex encoder. The core of our TGF strategy is the SIGs, developed based on Gated Fully Fusion (GFF) [54], which fuses features from multiple scales using gates that control the propagation of useful information. This enables the features at each scale to be enhanced by both deeper, semantically stronger features and shallower, spatially richer features, significantly reducing noises during feature fusion. Nonetheless, GFF is primarily regarded as a post-processing technique (multi-scale features are required to be yielded before being processed by GFF) and is infeasible to fuse heterogeneous features, which are extracted and fused progressively. In contrast, our developed SIGs (see Fig. 1) selectively inherit useful features in $\mathbf{X}_{i-1}^{G,F}$ from the previous layer into $\mathbf{X}_{i}^{G,F}$, the features at the current layer, where $i\in[1,n]\cap\mathbb{Z}$ denotes the layer number, and the superscripts ‘$G$’ as well as ‘$F$’ represent ‘geometric’ and ‘fused’ features [^1] It is worth noting here that we use $\mathbf{X}_{i}^{F}$ instead of $\mathbf{X}_{i}^{C}$, where the superscript ‘$C$’ denotes ‘contextual’ features, primarily because the branch processing the RGB images progressively fuses the geometric features extracted from the disparity maps to obtain the fused features., respectively. Our SIG outputs $\mathbf{\tilde{X}}_{i}^{G,F}$, selectively inherited feature maps at the $i$-th layer using the following expression:
@@
where $\Omega_{i}$ represents the SIG operation at the $i$-th layer, $\mathbf{1}_{k}$ denotes a column vector of ones, $\mathbf{G}_{i}\in[0,1]^{H_{i}\times W_{i}}$ represents a gate map that controls feature propagation, $\odot$ denotes the element-wise multiplication broadcasting in the channel dimension, and $\mathcal{R}$ represents the remapping operation, as detailed in [16].
@@
where $\mathbf{D}^{L}$ denotes the estimated disparity map, $\mathbf{F}^{C}_{i}$, $\mathbf{F}^{G}_{i}$, and $\mathbf{F}^{F}_{i}$ respectively represent the extracted contextual feature maps in the left view, geometric feature maps, and fused feature maps at the $i$-th layer, $\mathcal{E}^{G}_{i}$ and $\mathcal{E}^{F}_{i}$ denote the geometric and fused features encoding operations at the $i$-th layer, respectively, $\oplus$ represents the element-wise summation. Considering that the shallower features between the two tasks have similar numbers of channels, we make the contextual feature maps of the first three layers share weights with the feature maps extracted from the stereo matching network in our practical implementation. Our proposed TGF strategy selectively propagates features to subsequent layers, reducing indiscriminate heterogeneous feature fusion which can severely mislead semantic segmentation as the network goes deeper, thus achieving tighter coupling between these two relevant environmental perception tasks.

### v4

@@ Abstract
Semantic segmentation and stereo matching, respectively analogous to the ventral and dorsal streams in our human brain, are two key components of autonomous driving perception systems. Addressing these two tasks with separate networks is no longer the mainstream direction in developing computer vision algorithms, particularly with the recent advances in large vision models and embodied artificial intelligence. The trend is shifting towards combining them within a joint learning framework, especially emphasizing feature sharing between the two tasks. The major contributions of this study lie in comprehensively tightening the coupling between semantic segmentation and stereo matching. Specifically, this study makes three key contributions: (1) a tightly coupled, gated feature fusion strategy, (2) a hierarchical deep supervision strategy, and (3) a coupling tightening loss function. The combined use of these technical contributions results in TiCoSS, a state-of-the-art joint learning framework that simultaneously tackles semantic segmentation and stereo matching. Through extensive experiments on the KITTI, vKITTI2, and Cityscapes datasets, along with both qualitative and quantitative analyses, we validate the effectiveness of our developed strategies and loss function. Our approach demonstrates superior performance compared to prior arts, with a notable increase in mean intersection over union by over 9%.
Note to Practitioners
TiCoSS is a robust and effective joint learning framework that can simultaneously tackle semantic segmentation and stereo matching tasks. This work aims to improve semantic segmentation performance by exploring the potential complementarity and tightening the coupling between these two tasks. In the future, we plan to further improve the efficiency of the framework, so as to enable its real-time performance on resource-constrained hardware.
@@
semantic segmentation, stereo matching, autonomous driving, computer vision, joint learning.
@@ I-A Background
VISUAL environment perception serves as a fundamental and front-end module in robotic systems [1]. Semantic segmentation and stereo matching are two key visual environment perception tasks [2]. The former, akin to the ventral stream in our brain, provides a pixel-level understanding of the scene [3], while the latter, akin to the dorsal stream in our brain, mimics human binocular vision to acquire depth information [4], which is crucial for 3D geometry reconstruction. These two tasks collaborate to deliver both contextual and geometric information, resulting in a comprehensive scene understanding that significantly enhances the capabilities of robotic systems [5].
Nevertheless, previous studies [6, 7, 8] address these two tasks with separate networks, which limits their potential to share informative contextual and geometric information [9]. For instance, stereo matching networks can occasionally produce ambiguous disparity estimations, particularly in texture-less and occluded regions [10]. Semantic segmentation can provide pixel-level scene understanding results, which help resolve such ambiguities and ultimately lead to more reliable disparity estimations [11]. In addition, semantic segmentation networks often struggle to distinguish clear object boundaries, particularly in complex driving scenarios, due to the lack of spatial geometric information [12, 13]. A common solution for improved semantic segmentation performance is to employ feature-fusion networks equipped with duplex encoders to extract heterogeneous features from RGB-X data [14], where “X” provides spatial geometric information, such as the depth images generated from LiDAR point clouds [15] and surface normal maps obtained through depth-to-normal translation [3]. However, the availability and quality of “X” significantly influence the overall performance of semantic segmentation and can potentially limit the practical deployment of feature-fusion networks [5].
Therefore, in recent years, the simultaneous learning, deployment, and inference of both tasks have become a mainstream [16, 17, 18], because a unified joint learning framework can process contextual and geometric information more comprehensively. It also enables end-to-end training of the entire system, capable of tackling the challenges posed by both tasks [19, 20]. Consequently, this joint learning approach can enhance the overall performance of both semantic segmentation and stereo matching, and outperform models trained separately for each task [5].
@@ I-B Existing Challenges and Motivation
The performance of a feature-fusion semantic segmentation network is heavily influenced by the employed strategy for heterogeneous feature fusion [21, 22]. Currently, the bottleneck lies in the simplistic and indiscriminate fusion of heterogeneous features, which often causes conflicting learning representations and erroneous segmentation results [23]. Taking the state-of-the-art (SoTA) joint learning method S3M-Net [5] as an example, its adopted feature fusion strategy essentially performs an element-wise summation between contextual and geometric feature maps at each stage. These feature maps are then directly fed into subsequent layers without filtering out irrelevant information, leading to a loose coupling between the semantic segmentation and stereo matching tasks within the encoder. Furthermore, as the network goes deeper, such an indiscriminate feature fusion strategy tends to diminish the proportion of informative geometric features in the decoder’s input [24], potentially leading to unsatisfactory semantic segmentation performance.
Additionally, due to the vanishing gradient problem, existing joint learning frameworks often suffer from slow convergence during training. A common solution is to employ the deep supervision (DS) strategies that incorporate additional pathways to achieve gradient propagation [25]. Nonetheless, existing DS strategies employed in feature-fusion networks typically overlook the potential interactions between the main and side auxiliary classifiers, which can limit the overall semantic segmentation performance within our joint learning framework.
In the loss function aspect, previous joint learning frameworks, such as SegStereo [17] and SSNet [26], typically compute the losses for two tasks independently, and supervise the entire training process by simply minimizing the weighted sum of these losses. Such loss function fails to leverage the potential complementarity between the two tasks at the output level, which also results in a loose coupling.
Prior arts, such as S3M-Net [5] and DSNet [9], primarily focus on introducing a joint learning framework that performs semantic segmentation and stereo matching simultaneously. However, exploring the potential complementarity and tightening the coupling between these two tasks have received relatively limited attention in this research area and warrant further investigation.
@@ I-C Contributions
To address the aforementioned limitations, we introduce Tightly-Coupled Semantic Segmentation and Stereo Matching Network (TiCoSS), an end-to-end joint learning approach that focuses primarily on improving the coupling between stereo matching and semantic segmentation, which has not been emphasized in previous relevant studies. Our proposed TiCoSS introduces three new techniques: (1) a tightly-coupled, gated feature fusion (TGF) strategy, which utilizes a series of selective inheritance gates (SIGs) to propagate useful contextual and geometric information from the preceding layer to the current layer, resulting in a tightly-coupled encoder; (2) a hierarchical deep supervision (HDS) strategy that uses the fused feature maps with the highest resolution to guide deep supervision throughout each branch, as these features contain the most abundant local spatial details; (3) a novel coupling tightening (CT) loss, consisting of a widely used stereo matching loss presented in the study [8], the semantic consistency-guided (SCG) loss introduced in the study [5], a disparity inconsistency-aware (DIA) loss that leverages disparity estimation results to help distinguish clearer object boundaries, and a deep supervision consistency constraint (DSCC) loss which employs the Kullback-Leibler (KL) divergence to improve prediction consistency across outputs from all deep supervision branches. These contributions collectively advance S3M-Net, and results in TiCoSS, a new, powerful, and tightly-coupled joint learning framework that simultaneously performs robust and accurate semantic segmentation and stereo matching tasks. Extensive experiments conducted on the vKITTI2 [27], KITTI 2015 [28], and Cityscapes [29] datasets unequivocally demonstrate the effectiveness of the aforementioned contributions and the superior performance of TiCoSS over other SoTA approaches.
@@
  • The DIA loss and the DSCC loss that tighten the coupling between the two tasks, thereby further improving the semantic segmentation performance.
@@ II-A Semantic Segmentation
Semantic segmentation has been a long-standing challenge in the fields of computer vision and robotics over the past decade [1]. SoTA networks generally fall into two groups: (1) single-modal networks (with a single encoder) and (2) feature-fusion networks (with multiple encoders) [3]. Early efforts primarily focused on encoder-decoder architectures for pixel-level classification. Representative examples include the DeepLab series [30], as well as Transformer-based networks [31, 7, 32]. The encoder extracts hierarchical, contextual feature maps from input images, while the decoder generates segmentation maps by upsampling and combining feature maps from different encoder layers. Nonetheless, these networks are limited in effectively combining heterogeneous features extracted from different sources (or modalities) of visual information, which makes it challenging to produce accurate segmentation results in scenarios with poor lighting and illumination conditions [3]. This has led researchers to focus on feature-fusion networks that can effectively fuse heterogeneous features extracted from multiple sources (or modalities) of visual information. This problem is often referred to as “RGB-X semantic segmentation”, where “X” represents the additional modality (or source) of visual information, in addition to the RGB images. The most representative feature-fusion networks include convolutional neural network (CNN)-based ones, such as RTFNet [33] and the SNE-RoadSeg series [3, 25, 24], as well as Transformer-based ones, such as OFF-Net [34], RoadFormer [14], and DFormer [35]. In this article, we design TiCoSS based on S3M-Net, with a special emphasis on exploring more effective solutions for tighter coupling between semantic segmentation and stereo matching.
@@ II-B Stereo Matching
Owing to recent advancements in deep learning techniques, end-to-end deep stereo matching networks [36, 37, 8, 38] have dramatically outperformed traditional explicit programming-based stereo matching algorithms. PSMNet [36] introduces a spatial pyramid to capture multi-scale information and employs a series of 3D convolutional layers to aggregate both local and global contexts for cost computation. To address the high computational cost of 3D convolutions, researchers have sought ways to balance efficiency and accuracy in stereo matching. LEA-Stereo [37], for instance, introduces the first neural architecture search (NAS) framework to stereo matching, enabling automated architecture optimization. RAFT-Stereo [8], a rectified stereo matching approach, developed based on RAFT [39], uses a series of gated recurrent units to iteratively refine correlation features and improve disparity estimation accuracy. CRE-Stereo [38] further advances this approach by introducing an adaptive group-wise correlation layer to mitigate the impact of rectification errors in stereo images, resulting in more accurate disparity estimation results. In this article, we primarily focus on improving semantic segmentation performance, and therefore, adopt the stereo matching approach used in S3M-Net.
@@ II-C Simultaneous Semantic Segmentation and Stereo Matching
Existing joint learning frameworks that simultaneously address these two tasks mainly focus on improving disparity accuracy by leveraging semantic information [17, 16, 11, 18, 9]. However, discussions on utilizing disparity information to enhance semantic segmentation performance at the feature level for joint learning remain limited, except for the aforementioned “RGB-X semantic segmentation”. These prior arts often require extensive annotated training data or involve complex training strategies for joint learning. For example, SegStereo [17] necessitates an initial unsupervised training phase on the large-scale Cityscapes [29] dataset, followed by a subsequent supervised fine-tuning on the smaller KITTI [40, 28] datasets. Similarly, the approaches introduced in [11, 18, 41] require pre-training their spatial branches for stereo matching on the large-scale SceneFlow [42] dataset before fine-tuning both semantic and spatial branches on the KITTI [28] dataset. DSNet [9] adopts a different joint learning strategy by alternating the training of the semantic segmentation and stereo matching networks, with parameters of each network being frozen during the training of the other. Nevertheless, leveraging both contextual and geometric information can be challenging, as the shared features between these two tasks are not learned in an end-to-end manner. DispSegNet [16] utilizes an embedding learned from the semantic segmentation branch to refine the initial disparity estimations. RTS2Net [18] relies on coarse-to-fine estimations in a multi-stage fashion for accurate disparity estimation. However, both methods only achieve limited improvement in semantic segmentation since they fail to leverage informative spatial details to enhance semantic segmentation performance. SSNet [26] employs a single encoder to extract shareable features for both tasks. However, as demonstrated in the study [43], such shareable features may not be suitable for both dense prediction and geometric vision tasks. S3M-Net [5], MENet [12], and SG-RoadSeg [19] use two separate encoding branches to accomplish these two tasks simultaneously, but the weak coupling between these branches limits the integration of contextual and geometric information. In contrast to the aforementioned approaches, our proposed TiCoSS uses a tightly-coupled joint learning framework that effectively leverages both contextual and geometric information. Moreover, TiCoSS is trained in an end-to-end manner and is capable of learning accurate and robust semantic segmentation and stereo matching tasks simultaneously, even with limited training data.
@@ III-B Tightly-coupled Gated Fusion Strategy
S3M-Net [5] proposes an effective joint learning framework to simultaneously perform semantic segmentation and stereo matching. Despite achieving impressive results, these two tasks are loosely coupled. It merely employs the feature fusion strategy proposed in SNE-RoadSeg [3], where the geometric features extracted from disparity maps are indiscriminately fused into the contextual features extracted from RGB images via simplistic element-wise summation. The fused heterogeneous features are then treated as preceding contextual features and fed into subsequent layers without selective processing, which can potentially mislead the semantic segmentation task. This is primarily because the deeper geometric features contain irrelevant semantic information, and as the network goes deeper, the proportion of contextual features in the decoder’s input tends to diminish [24].
Our TGF strategy is, therefore, designed to overcome this limitation by selectively complementing contextual features with informative geometric features, resulting in a tightly-coupled duplex encoder. The core of our TGF strategy is the SIGs, developed based on Gated Fully Fusion (GFF) [44], which fuse features from multiple scales using gates that control the propagation of useful information. This enables the features at each scale to be enhanced by both deeper, semantically stronger features and shallower, spatially richer features, significantly reducing noises during feature fusion. Nonetheless, GFF is primarily regarded as a late fusion strategy [45], as it performs feature fusion at the decision layer and requires multi-scale features to be generated prior to processing. Additionally, GFF focuses solely on fusing features extracted from RGB images across multiple scales and is not well-suited to fuse heterogeneous features which are extracted and fused progressively. In contrast, our proposed TGF strategy performs intermediate feature fusion during the encoding stage, enabling more interactions between heterogeneous features. Specifically, it utilizes a series of SIGs (see Fig. 1) to selectively inherit useful information in $\mathbf{X}_{i-1}^{G,F}$ from the previous layer into $\mathbf{X}_{i}^{G,F}$, the features at the current layer, where $i\in[1,n]\cap\mathbb{Z}$ denotes the layer number, and the superscripts ‘$G$’ as well as ‘$F$’ represent ‘geometric’ and ‘fused’ features [^1] It is worth noting here that we use $\mathbf{X}_{i}^{F}$ instead of $\mathbf{X}_{i}^{C}$, where the superscript ‘$C$’ denotes ‘contextual’ features, primarily because the branch processing RGB images progressively fuses the geometric features extracted from disparity maps to obtain fused features., respectively. Our SIG outputs $\mathbf{\tilde{X}}_{i}^{G,F}$ selectively inherit feature maps at the $i$-th layer using the following expression:
@@
where $\Omega_{i}$ represents the SIG operation at the $i$-th layer, $\mathbf{1}_{k}$ denotes a column vector of ones, $\mathbf{G}_{i}\in[0,1]^{H_{i}\times W_{i}}$ represents a gate map that controls feature propagation, $\odot$ denotes the element-wise multiplication broadcasting in the channel dimension, and $\mathcal{R}$ represents the remapping operation, as detailed in the study [5].
@@
where $\mathbf{D}^{L}$ denotes the estimated disparity map, $\mathbf{F}^{C}_{i}$, $\mathbf{F}^{G}_{i}$, and $\mathbf{F}^{F}_{i}$ respectively represent the extracted contextual feature maps in the left view, geometric feature maps, and fused feature maps at the $i$-th layer, $\mathcal{E}^{G}_{i}$ and $\mathcal{E}^{F}_{i}$ denote the geometric and fused features encoding operations at the $i$-th layer, respectively, and $\oplus$ represents the element-wise summation. Considering that the shallower features between the two tasks have similar numbers of channels, we make the contextual feature maps of the first three layers share weights with the feature maps extracted from the stereo matching network in our practical implementation. Our proposed TGF strategy selectively propagates useful information to subsequent layers, reducing indiscriminate heterogeneous feature fusion which can severely mislead semantic segmentation as the network goes deeper, thus achieving tighter coupling between these two relevant perception tasks. Further theoretical analyses of the TGF strategy are provided in the supplement.